OpenStudy (anonymous):

find the x- and y-intercepts of the line 3x – 12y = 6.

OpenStudy (lgbasallote):

y intercept -> set x = 0; find y x intercept - > set y= 0; find x..willl you try?? :DDD

OpenStudy (anonymous):

jst plug in each x, y =0 for each

OpenStudy (igloo3000):

plug in 0 for x if you are trying to find y and put in 0 for y if you want to find x

OpenStudy (anonymous):

set x=0 for y-intercept and set y=0 for x-intercept!!

OpenStudy (anonymous):

SO for y intercept you get: 3(0)-12y=6 -12y=6 y=-1/2

OpenStudy (anonymous):

for x-intercept you get: 3x-12(0)=6 3x=6 x=2